### v1

@@ 5 Comparison with recent literature results
In this section, we compare our results with those found in the literature in recent years. Similar to what was done in T23, we focus our attention on studies involving either Cepheids or Open Clusters (OCs) since they span a wide range of ages and represent optimal tracers. Since our results are in fair agreement with our previous work, we refer to T23 for a comparison with less recent literature works to avoid redundancies in the discussion.
Recently, da Silva et al. (2023) collected 356 DCEPs and compared the linear and logarithmic regression along the Galactocentric distance axis for Fe, O and S, finding a change of slope which flattens after around 12.5 kpc, a feature not found in this work. Regarding the linear fit, their values strongly disagree with ours:
@@
Fe:$(-0.041\pm 0.003)Rg+(0.32\pm 0.02)$
As already stated in their work, a possible explanation could be related to the different models used to derive abundances (see their section 4.1) and distances (see their section 2.3). Another cause of this discrepancy could be related to the different amount of targets available at $R_{GC}$¿15kpc. Nonetheless, similarly to what was stated in their work, we find a strong difference between the O and S gradients, the latter being considerably steeper than the former.
@@
A similar discussion applies to the study by Joshi et al. (2024) who collected a remarkable sample of 1879 Ocs to study the evolution of metallicity in the Galactic disk. About 90% of their objects are younger than 1 Gyr and the resulting slope (for $R_{GC}$¡12.8kpc) is in perfect agreement with our work ($-0.070\pm 0.002$ dex $kpc^{-1}$).
In their investigation on the origin of neutron-capture elements, Molero et al. (2023) derived chemical evolution models starting from the two-infall model by Chiappini et al. (1997) and confronted with observational data from the 6th Data Release of the $Gaia-$ESO survey. Their observation (with OCs) and prediction (Model R-150) for the present time radial gradient for [Fe/H] and [Eu/H] is in good agreement with our results (see their table 2).
Except for a slightly shallower slope documented in Netopil et al. (2022) ($-0.058$ dex kpc-1), derived from a merged dataset of individual studies and the APOGEE (Apache Point Observatory Galactic Evolution Experiment) survey comprising 136 OCs, our findings align well with several other OC studies previously explored in depth by Magrini et al. (2023), specifically those by Spina et al. (2022), Zhang et al. (2021), and Myers et al. (2022), which reported slopes of $-0.064\pm 0.007$, $-0.066\pm 0.005$, and $-0.073\pm 0.002$ dex kpc-1, respectively.
On the contrary, Myers et al. (2022) identified a somewhat steeper slope, although statistically consistent when focusing solely on young OCs (age ¡ 1Gyr) (refer to table 2 of Magrini et al. (2023)). Noteworthy are the findings of Magrini et al. (2023), particularly in their tables A.8 and A.10, where linear weighted fits were conducted for various chemical species. Overall, strong concordance is observed both in the overall fit and when restricting the analysis to young OCs (age ¡ 1Gyr).
@@ 6 Summary
In the context of the C-MetaLL project, we presented the results obtained from 331 high-resolution spectra collected with the UVES@VLT, HARPS-N@TNG and ESPADONS@CFHT instruments for a sample of 180 individual DCEPs located over a wide range of Galactocentric radii (5$<R_{GC}<20$ kpc).
@@
For each target, we derived accurate atmospheric parameters, radial velocities, and abundances for up to 29 different species. The iron abundances range between +0.5 and --1 dex with a rather even distribution of DCEPs with metallicity. The sample presented in this paper was complemented with the data already published in the context of the C-MetaLL survey, resulting in a total of 292 pulsators whose spectra have been analysed in a homogeneous way.
@@
The main aim of these observations is the derivation of the metallicity dependence of DCEP multi-wavelength $PL$ relations. This subject will be addressed in a forthcoming paper. In this work, we exploited our homogeneous sample to study the abundance gradients of the Galactic disc in a range of Galactocentric radii spanning the range of 5-20 kpc.
Here we report the main results of this work:
@@
  • Studying the relation between each chemical species (in their [X/Fe] form) and iron we observed an $\alpha$ enhancement for [Fe/H]<absent<-< -0.5 dex). This is a well-known occurrence, already observed both empirically (Hayden et al., 2015; Duffau et al., 2017; Trentin et al., 2023) and theoretically (Tinsley, 1979; Matteucci & Brocato, 1990; Palicio et al., 2023). At higher metallicities, this negative trend flattens.
@@
  • Iron peak and heavy elements generally present a negative gradient compared with the iron content, while those produced via s-process (namely Zr, Ba, La) show an arch-form behaviour (Molero et al., 2023), with a change of trend at [Fe/H] $\approx$--0.2 dex.
@@
  • For some elements (Sc, Cu, Zn and Zr), we found an apparent bias in the abundances between our previous work and this one at lower metallicities, with Zn having an opposite behaviour compared with what appears for Sc, Cu and Zr. Based on a direct visual comparison of the spectra for stars with similar atmospheric parameters and iron content, we confirmed the difference in the line depths for the available spectral lines for these elements.
@@
  • As expected, we found a clear negative radial gradient for most of the elements (except Ce, Pr, Sm and Gd). Our targets are evenly distributed over a broad range of distances. When complementing our sample with other literature stars from other works (as done in T23 for example), the slopes do not significantly change. This confirms the importance of using a homogeneous dataset. Good agreement is found with other DCEP and OC works in the most recent literature.
@@
  • Focusing on the case of Iron, we did not find any evidence for a break at any galactic radii and the estimated slope results are more negative ($-0.071\pm 0.003$ dex kpc-1) compared with the single-line case of our previous work. When considering the two-line case, this value is in between the inner and outer slopes.
@@
  • Overimposing the polar distribution of our dataset with the Galactic spiral arms (Reid et al., 2019; Minniti et al., 2021), we found a qualitative fit for our farthest star, which traces both the Outer and the extension of OSC arms. Should this result be confirmed in further works, it would be possible to put better constraints for both the spiral structures in a region still poorly explored (in particular for the OSC arm).
@@
  • The possible association of our targets to the two spiral arms mentioned above could qualitatively justify the apparent change of abundance trend for Sc, Cu, Zn and Zr. Interestingly enough, the polar representation of the distribution of these four elements shows a peculiar behaviour of zinc. This species appears more abundant in the stars associated with the Outer arm, while the opposite is visible for Sc, Cu and Zr. No clear explanation has been found for this behaviour.
@@
Finally, the homogeneous sample of DCEPs with metallicities from high-resolution spectroscopy will be exploited in a forthcoming paper for the study of the metallicity dependence of the DCEP $PL$ relations.
@@
(dex $kpc^{-1}$) (dex) (dex)
C -0.066 $\pm$ 0.003 0.324 $\pm$ 0.035 0.13
O -0.046 $\pm$ 0.004 0.352 $\pm$ 0.048 0.18
Na -0.066 $\pm$ 0.003 0.876 $\pm$ 0.040 0.16
Mg -0.058 $\pm$ 0.004 0.531 $\pm$ 0.047 0.19
Al -0.048 $\pm$ 0.003 0.531 $\pm$ 0.034 0.11
Si -0.051 $\pm$ 0.002 0.525 $\pm$ 0.028 0.06
S -0.062 $\pm$ 0.003 0.641 $\pm$ 0.031 0.12
Ca -0.059 $\pm$ 0.003 0.475 $\pm$ 0.034 0.14
Sc -0.068 $\pm$ 0.006 0.810 $\pm$ 0.070 0.33
Ti -0.054 $\pm$ 0.003 0.513 $\pm$ 0.041 0.15
V -0.030 $\pm$ 0.003 0.285 $\pm$ 0.039 0.15
Cr -0.052 $\pm$ 0.003 0.482 $\pm$ 0.037 0.13
Mn -0.075 $\pm$ 0.004 0.497 $\pm$ 0.042 0.16
Fe -0.071 $\pm$ 0.003 0.588 $\pm$ 0.029 0.12
Co -0.039 $\pm$ 0.004 0.464 $\pm$ 0.051 0.20
Ni -0.064 $\pm$ 0.003 0.509 $\pm$ 0.033 0.10
Cu -0.039 $\pm$ 0.005 0.281 $\pm$ 0.058 0.25
Zn -0.070 $\pm$ 0.003 0.458 $\pm$ 0.034 0.09
Sr -0.032 $\pm$ 0.010 0.494 $\pm$ 0.091 0.15
Y -0.046 $\pm$ 0.004 0.417 $\pm$ 0.048 0.21
Zr -0.036 $\pm$ 0.004 0.526 $\pm$ 0.056 0.22
Ba -0.046 $\pm$ 0.006 0.461 $\pm$ 0.073 0.33
La -0.033 $\pm$ 0.003 0.356 $\pm$ 0.043 0.18
Ce -0.003 $\pm$ 0.002 0.040 $\pm$ 0.026 0.10
Pr 0.002 $\pm$ 0.004 -0.325 $\pm$ 0.044 0.13
Nd -0.033 $\pm$ 0.003 0.273 $\pm$ 0.039 0.16
Sm -0.005 $\pm$ 0.008 0.054 $\pm$ 0.079 0.15
Eu -0.035 $\pm$ 0.004 0.246 $\pm$ 0.048 0.12
Gd 0.008 $\pm$ 0.006 0.124 $\pm$ 0.069 0.22
 [^9] Coefficients of the linear fit of the form $[X/H]=\alpha\times R_{gc}+\beta$ with relative dispersion and Pearson coefficient.
@@ Acknowledgements.
This research has made use of the SIMBAD database operated at CDS, Strasbourg, France. We acknowledge funding from INAF GO-GTO grant 2023 “C-MetaLL - Cepheid metallicity in the Leavitt law” (P.I. V. Ripepi). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix A Stars with uncertain classification
As mentioned in Sect. 2.1, the three stars Gaia_DR2_4087335043492541696, ASAS-SN_J165340.10-332041.7 and V532_Sco deserve an in-depth analysis to ascertain whether or not they could be type II Cepheids. To this aim, we plot in Fig. 13 the light curves in the $Gaia$ bands for the three stars, where the phases at which the spectra were observed are reported. These light curves were used by Ripepi et al. (2023) to classify these objects as fundamental DCEPs. However, the three stars have been classified as belonging to the W Vir class (type II Cepheids) by several authors in the literature. In more detail:
@@
  • Gaia_DR2_4087335043492541696, also known as V410 Sgr or OGLE BLG-T2CEP-1340 has been classified as W Vir by several authors in the literature, e.g. Jurkovic et al. (2023) based on Kepler K2 data and Soszyński et al. (2020) based on $V,I$ photometry in the context of the OGLE (Optical Gravitational Lensing Experiment) survey, just to mention the most recent works.
@@
  • ASAS-SN_J165340.10-332041.7, also known as OGLE BLG-T2CEP-1089 has been considered as DCEP by Jayasinghe et al. (2018) and Skowron et al. (2019), but lately re-classified as W Vir by Soszyński et al. (2020).
@@
  • V532_Sco, also known as SV HV 10484 or OGLE BLG-T2CEP-1229 has been classified as W Vir by Harris (1985), a classification subsequently confirmed by Soszyński et al. (2020).
@@
We analyse now the spectra of the three stars searching for features which could possibly allow us to classify them in terms of DCEPs or W Vir stars. As showed in Fig. 14, we focus our attention to the $H_{\alpha}$ region where, according to Schmidt et al. (2004) two features, due to the presence of violent shocks in the moving atmospheres of the type II Cepheids can help us to distinguish between Type I and type II Cepheids: i) the difference in velocity between the $H_{\alpha}$ and the metallic lines and ii) the presence of strong emission in $H_{\alpha}$ from approximately half the cycle till phase 0.1 after the maximum. The left panel of Fig. 14 shows the three epoch spectra of Gaia_DR2_4087335043492541696 and the single-epoch spectra for ASAS-SN_J165340.10-332041.7 and V532_Sco. For comparison, the right panel shows the spectra of different DCEPs with periods and phases similar to those of the three investigated stars. The difference in velocity between $H_{\alpha}$ and the metallic lines can be clearly seen for Gaia_DR2_4087335043492541696, especially at ϕ=italic-$\phi=$0.13 and 0.51, and for ASAS-SN_J165340.10-332041.7. However, similar differential velocities can also be seen in DCEPs, therefore, as noted by Schmidt et al. (2004), this feature is not a good discriminant, at least for periods larger than 8 days. However, the presence of $H_{\alpha}$ emission or P-Cygni profiles (both direct and inverse) is evident in almost all the left-panel spectra, while it is almost absent among DCEPs. An exception appears to be V532_Sco, whose spectrum is the most similar to the relative DCEPs’.
@@
On these bases we conclude that Gaia_DR2_4087335043492541696 and ASAS-SN_J165340.10-332041.7 are almost certainly W Vir pulsators, while the classification of V532_Sco remains uncertain.

### v3

@@ 5 Comparison with recent literature results
In this section, we compare our results with those found in the literature in recent years. Similar to what was done in T23, we focus our attention on studies involving either Cepheids or OCs, since they span a wide range of ages and represent optimal tracers. Since our slope is in agreement with our previous work, we refer to this paper for a comparison with less recent literature works to avoid redundancies in the discussion.
Recently, da Silva et al. (2023) collected 356 DCEPs and compared the linear and logarithmic regression along the Galactocentric distance axis for [Fe/H], [O/H], and [S/H], finding a change of slope that flattens after around 12.5 kpc, a feature not found in this work. Regarding the linear fit, their values strongly disagree with ours:
@@
Fe:$(-0.041\pm 0.003)R_{GC}+(0.32\pm 0.02)$
As was already stated in their work, a possible explanation could be related to the different models used to derive abundances (see their section 4.1) and distances (see their section 2.3). Another cause of this discrepancy could be related to the different amount of targets available at $R_{GC}$¿15kpc. Nonetheless, similarly to what was stated in their work, we find a strong difference between the O and S gradients, the latter being considerably steeper than the former.
@@
A similar discussion applies to the study by Joshi et al. (2024), who collected a remarkable sample of 1879 OCs to study the evolution of metallicity in the Galactic disc. About 90% of their objects are younger than 1 Gyr and the resulting slope (for $R_{GC}$¡12.8kpc) is in agreement with our work ($-0.070\pm 0.002$ dex $kpc^{-1}$) at 1.5 $\sigma$.
In their investigation into the origin of neutron-capture elements, Molero et al. (2023) derived chemical evolution models, starting with the two-infall model by Chiappini et al. (1997), and confronted them with observational data from the 6th Data Release of the $Gaia-$ESO survey. Their observation (with OCs) and prediction (Model R-150) for the present-time radial gradient for [Fe/H] and [Eu/H] are in good agreement with our results (see their table 2).
Except for a slightly shallower slope documented in Netopil et al. (2022) ($-0.058$ dex kpc-1), derived from a merged dataset of individual studies and the Apache Point Observatory Galactic Evolution Experiment (APOGEE) survey comprising 136 OCs, our findings align well with several other OC studies previously explored in depth by Magrini et al. (2023), specifically those by Spina et al. (2022) and Zhang et al. (2021), which report slopes of $-0.064\pm 0.007$ and $-0.066\pm 0.005$, respectively.
On the contrary, Myers et al. (2022) identify a somewhat steeper slope ($-0.073\pm 0.002$ dex kpc-1), although one that is statistically consistent when focussing solely on young OCs (age ¡ 1Gyr) (refer to table 2 of Magrini et al. (2023)). Of note are the findings of Magrini et al. (2023), particularly in their tables A.8 and A.10, for which linear weighted fits were conducted for various chemical species. Overall, strong concordance is observed both in the overall fit and when restricting the analysis to young OCs (age ¡ 1Gyr).
@@ 6 Summary
In the context of the C-MetaLL project, we have presented the results obtained from 331 HiRes spectra collected with the UVES/VLT, HARPS-N/TNG, and ESPaDOnS/CFHT instruments for a sample of 180 individual DCEPs located over a wide range of Galactocentric radii (5$<R_{GC}<20$ kpc). For each target, we derived accurate atmospheric parameters, radial velocities, and abundances for up to 29 different species. The iron abundances range between +0.5 and --1 dex with a rather even distribution of DCEPs with metallicity. The sample presented in this paper was complemented with the data already published in the context of the C-MetaLL survey, resulting in a total of 292 pulsators whose spectra have been analysed in a homogeneous way.
@@
In this work, we have exploited our homogeneous sample to study the abundance gradients of the Galactic disc in a range of Galactocentric radii spanning the range of 5-20 kpc. Here, we report the main results of this work:
@@
  • Studying the relation between each chemical species (in its [X/Fe] form) and iron, we observe an $\alpha$ enhancement for [Fe/H]<absent<-< -0.5 dex. This is a well-known occurrence, already observed both empirically (Hayden et al., 2015; Duffau et al., 2017; Trentin et al., 2023) and theoretically (Tinsley, 1979; Matteucci & Brocato, 1990; Palicio et al., 2023). At higher metallicities, this negative trend flattens.
@@
  • Iron peak and heavy elements generally present a negative gradient compared with the iron content, while those produced via the s-process (namely, Zr, Ba, and La) show an arch-form behaviour (Molero et al., 2023), with a change of trend at [Fe/H] $\approx$--0.2 dex.
@@
  • For some elements (Sc, Cu, Zn, and Zr), we have found an apparent bias in the abundances between T23 and this study at lower metallicities, with Zn having an opposite behaviour to what appears for Sc, Cu, and Zr. Based on a direct visual comparison of the spectra for stars with similar atmospheric parameters and iron content, we have confirmed the difference in the line depths for the available spectral lines for these elements.
@@
  • As was expected, we have found a clear negative radial gradient for most of the elements (except Ce, Pr, Sm, and Gd). Our targets are evenly distributed over a broad range of distances. When complementing our sample with other literature stars from other works (as was done in T23, for example), the slopes do not significantly change. Good agreement is found with other DCEP and OC works in the most recent literature.
@@
  • Focussing on the case of iron, we did not find any evidence for a break at any galactic radii and the estimated slope ($-0.064\pm 0.003$ dex kpc-1) is in good agreement compared with both the single-line case of Trentin et al. (2023) and the inner slope (when considering their two-line case).
@@
  • Superimposing the polar distribution of our dataset with the Galactic spiral arms (Reid et al., 2019; Minniti et al., 2021), we found a qualitative fit for our farthest star, which traces both the Outer and the extension of OSC arms. Should this result be confirmed in further works, it would be possible to put better constraints on both the spiral structures in a region that is still poorly explored (in particular on the OSC arm).
@@
  • The possible association of our targets with the two spiral arms mentioned above could qualitatively justify the apparent change of abundance trend for Sc, Cu, Zn, and Zr. Interestingly enough, the polar representation of the distribution of these four elements shows a peculiar behaviour of zinc. This species appears more abundant in the stars associated with the Outer arm, while the opposite is visible for Sc, Cu, and Zr. No clear explanation has been found for this behaviour.
@@
Finally, the homogeneous sample of DCEPs with metallicities from HiRes spectroscopy will be exploited in a forthcoming paper for the study of the metallicity dependence of the DCEP $PL$ relations.
@@
(dex kpc-1) (dex) (dex)
C -0.067 $\pm$ 0.003 0.320 $\pm$ 0.035 0.13
O -0.044 $\pm$ 0.004 0.329 $\pm$ 0.050 0.19
Na -0.061 $\pm$ 0.003 0.828 $\pm$ 0.039 0.16
Mg -0.055 $\pm$ 0.004 0.488 $\pm$ 0.047 0.19
Al -0.044 $\pm$ 0.003 0.486 $\pm$ 0.035 0.12
Si -0.049 $\pm$ 0.002 0.501 $\pm$ 0.029 0.08
S -0.059 $\pm$ 0.003 0.596 $\pm$ 0.032 0.13
Ca -0.055 $\pm$ 0.003 0.434 $\pm$ 0.035 0.15
Sc -0.061 $\pm$ 0.006 0.736 $\pm$ 0.071 0.33
Ti -0.052 $\pm$ 0.004 0.488 $\pm$ 0.043 0.16
V -0.028 $\pm$ 0.003 0.259 $\pm$ 0.040 0.16
Cr -0.051 $\pm$ 0.003 0.458 $\pm$ 0.038 0.14
Mn -0.078 $\pm$ 0.004 0.529 $\pm$ 0.044 0.17
Fe -0.064 $\pm$ 0.003 0.530 $\pm$ 0.029 0.11
Co -0.035 $\pm$ 0.004 0.422 $\pm$ 0.053 0.20
Ni -0.061 $\pm$ 0.003 0.478 $\pm$ 0.034 0.11
Cu -0.035 $\pm$ 0.005 0.236 $\pm$ 0.059 0.26
Zn -0.070 $\pm$ 0.003 0.451 $\pm$ 0.035 0.11
Sr -0.032 $\pm$ 0.009 0.496 $\pm$ 0.088 0.15
Y -0.041 $\pm$ 0.004 0.366 $\pm$ 0.049 0.22
Zr -0.031 $\pm$ 0.005 0.474 $\pm$ 0.057 0.23
Ba -0.043 $\pm$ 0.006 0.429 $\pm$ 0.073 0.33
La -0.030 $\pm$ 0.004 0.319 $\pm$ 0.044 0.19
Ce -0.002 $\pm$ 0.002 0.037 $\pm$ 0.026 0.10
Pr 0.003 $\pm$ 0.004 -0.329 $\pm$ 0.044 0.13
Nd -0.031 $\pm$ 0.003 0.245 $\pm$ 0.039 0.16
Sm -0.008 $\pm$ 0.008 0.077 $\pm$ 0.077 0.15
Eu -0.032 $\pm$ 0.005 0.209 $\pm$ 0.054 0.13
Gd 0.003 $\pm$ 0.006 0.178 $\pm$ 0.071 0.21
 [^10] Coefficients of the linear fit of the form [X/H]=$\alpha\times R_{GC}+\beta$ with relative dispersion coefficient.
@@ Acknowledgements.
This research has made use of the SIMBAD database operated at CDS, Strasbourg, France. We acknowledge funding from INAF GO-GTO grant 2023 “C-MetaLL - Cepheid metallicity in the Leavitt law” (P.I. V. Ripepi). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. This research was supported by the International Space Science Institute (ISSI) in Bern/Beijing through ISSI/ISSI-BJ International Team project ID #24-603 - “EXPANDING Universe” (EXploiting Precision AstroNomical Distance INdicators in the Gaia Universe).
@@ Appendix A Stars with uncertain classification
As mentioned in Sect. 2.1, the three stars Gaia DR2 4087335043492541696, ASAS-SN J165340.10-332041.7 and V532 Sco deserve an in-depth analysis to ascertain whether or not they could be type II Cepheids. To this aim, we plot in Fig. 14 the light curves in the $Gaia$ bands for the three stars, where the phases at which the spectra were observed are reported. These light curves were used by Ripepi et al. (2023) to classify these objects as fundamental DCEPs. However, the three stars have been classified as belonging to the W Vir class (type II Cepheids) by several authors in the literature. In more detail:
@@
  • Gaia DR2 4087335043492541696, also known as V410 Sgr or OGLE BLG-T2CEP-1340 has been classified as W Vir by several authors in the literature, e.g. Jurkovic et al. (2023) based on Kepler K2 data and Soszyński et al. (2020) based on $V,I$ photometry in the context of the OGLE (Optical Gravitational Lensing Experiment) survey, just to mention the most recent works.
@@
  • ASAS-SN J165340.10-332041.7, also known as OGLE BLG-T2CEP-1089 has been considered as DCEP by Jayasinghe et al. (2018) and Skowron et al. (2019), but lately re-classified as W Vir by Soszyński et al. (2020).
@@
  • V532 Sco, also known as SV HV 10484 or OGLE BLG-T2CEP-1229 has been classified as W Vir by Harris (1985), a classification subsequently confirmed by Soszyński et al. (2020).
@@
We analyse now the spectra of the three stars searching for features which could possibly allow us to classify them in terms of DCEPs or W Vir stars. As showed in Fig. 15, we focus our attention to the $H_{\alpha}$ region where, according to Schmidt et al. (2004) two features, due to the presence of violent shocks in the moving atmospheres of the type II Cepheids can help us to distinguish between Type I and type II Cepheids: i) the difference in velocity between the $H_{\alpha}$ and the metallic lines and ii) the presence of strong emission in $H_{\alpha}$ from approximately half the cycle till phase 0.1 after the maximum. The left panel of Fig. 15 shows the three epoch spectra of Gaia DR2 4087335043492541696 and the single-epoch spectra for ASAS-SN J165340.10-332041.7 and V532 Sco. For comparison, the right panel shows the spectra of different DCEPs with periods and phases similar to those of the three investigated stars. The difference in velocity between $H_{\alpha}$ and the metallic lines can be clearly seen for Gaia DR2 4087335043492541696, especially at ϕ=italic-$\phi=$0.13 and 0.51, and for ASAS-SN J165340.10-332041.7. However, similar differential velocities can also be seen in DCEPs, therefore, as noted by Schmidt et al. (2004), this feature is not a good discriminant, at least for periods larger than 8 days. However, the presence of $H_{\alpha}$ emission or P-Cygni profiles (both direct and inverse) is evident in almost all the left-panel spectra, while it is almost absent among DCEPs. An exception appears to be V532 Sco, whose spectrum is the most similar to the relative DCEPs’.
@@
On these bases we conclude that Gaia DR2 4087335043492541696 and ASAS-SN J165340.10-332041.7 are almost certainly W Vir pulsators, while the classification of V532 Sco remains uncertain.
@@
Appendix B Phases for seven DCEPs devoid of $Gaia$ periods and epochs of maximum light.
To use epochs close to the spectroscopy observations we proceeded as follows:
  • AP Sgr: the period and epoch of maximum light was calculated directly from $Gaia$ DR3 $G$-band time series using the Period04 package (Lenz & Breger 2005). As a result, we obtained P=5.05790 days and Epoch=58992.77858 days, where the errors are on the last digits.
  • T Mon, U Sgr: for these stars, we have recent periods (at epoch 59591.5 days) provided by Csörnyei et al. (2022). To determine epochs close to our observations we used again Period04 but imposed the period instead of recalculating it. Therefore the adopted periods and epochs of maximum light are P=27.033911203 days; Epoch=59576.79823 days; P=6.745265182 days Epoch=59587.19128 days for T Mon and U Sgr, respectively.
  • ASAS J060722+0834.0, DP Mon, NSVS 2150508; V981 Mon: for these stars we adopted periods and epochs of maximum light from the All-Sky Automated Survey for Supernovae (ASAS-SN Shappee et al. 2014; Christy et al. 2023).